A short-term insurance company receives, in each working day, three motor vehicle claims. Assume that the daily claims rate follow a Poisson process. How likely is it that either eight or nine motor vehicle claims will be received in two working days? Express your answer to 4 decimal places.


1
Expert's answer
2021-07-05T14:28:03-0400

Let "X=" the number of received motor vehicle claims: "X\sim Po(\lambda t)."  

Given "\lambda t=3(2)=6"


"P(X=8\ or X=9)=\dfrac{e^{-6}\cdot6^8}{8!}+\dfrac{e^{-6}\cdot6^9}{9!}"

"\approx0.1721"